Run Coordination Meetings: Fall 2014
Daily Meetings, 8:45am, Hall D conference room. Held by the Run Coordinator. Present Run Coordinator: Alexandre Deur
To connect from the outside by phone please do the following
2.) To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 660743227.
3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.
4.) The moderator code is 8394
Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada
Specific instructions for connecting from JLab's Hall D conference room:
- Turn polycon on if necessary (do it before turning the computer on)
- With the polycon, place a call at 126.96.36.199
- Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
- You may have to unmute the microphone: #*4
- Turn the computer on if needed
Meeting General Agenda
Oct 30th - Dec 24th 2014 meetings
- Report relevant updates from 7:50 and 8:00 daily MCC meetings:
- Accelerator schedule: Accelerator is down due a JT valve problem in the south linac.
- Link to accelerator short term schedule
- Report on Work in the Hall:
- If you have work to do in the Hall, please document it in the electronic logbook (either update this entry or link to it) and notify me (RC) and Tom (Hall work coordinator).
- The upstream platform is now a hearing protection area. Hall D is a radiation aread. OSL need to be worn everywhere in Hall D.
Hot Check OUT Status 100% ready ; 00% checked ; 00% not ready
- Current runplan
- Commissioning readiness
- Installation punch list (items in red: must by Oct 15th ; Blue: Done by 2nd Fall run; Green: Nice to have it
The wire is installed. Tom measured the wire diameter without insulation to be 200 microns.
Chamber is closed and under vacuum, leak-checked and ready to go.
Pair spectrometer detector - Positioned on the stand. Survey is done. Profile monitors -
The two profilers are now installed and surveyed.
Active collimator -
Cables are terminated and in the Hall. All hardware work is completed.
The light leaks of the SC are largely fixed. SC is mounted back.
Tagger shielding - Wall (final labyrinth configuration) is completed and both small walls are done. Work on iron shielding near the photon photon dump is done. Pair Spectrometer Magnet remote control - Control is now working. Pair Spectrometer Magnet Field readout Done Survey:
Pair Spectrometer detector: finishing today.
Upstream Profiler in collimator cave.
Start counter: Survey should be done by the end of the day.
FCal: Still working on it. Should be completed today
CDC: Should be completed (confirmation?). Microscope
Beam pipe section
Vacuum - Done. Amorphous Radiators - FSD done and tested. Halo Counters - All are installed. The 8 in the Hall are successfully readout. The 7 in the tagger Hall are not readout and won't be available for the first part of the commissioning. Microscope - Microscope is in its correct position. Tagger crates - Now removed from the Hall. Hodoscope checks - Detectors and electronics were tested. 9 out of 170 counters have problems (1 not responsive, 8 with cabling problems) . Franz took them back to Catholic Univ. to fix them.
Radiation monitors - Gamma and neutron probes ready. Ion Chambers installed. Witness SiPM will be installed today.
2nd sweep magnet. (2nd collimator aperture) Working (tested at 100A. Nominal current: 197A). Not required for Fall run. </span>
- Specific safety notice (from Tom/Tim/Mark):
Commissioning of Each Detector
- FADC vs. CAEN TDC timing, November 21, 2014
- Overview of Offline Monitoring Nov 20 2014
- Start Counter ST performance during Run 995
- CDC CDC Update
- FCAL FCAL Update
- BCAL BCAL status from HDLOG
- TOF Status: TOF performance in Run 995
- Active collimator status: Initial calibration from y beam scan
- Offline Monitoring: Run Browser, Plot Browser, and Time Series webpages
Detector Daily Report
- 12/03/14: BCal report (Elton and Mark Dalton) 1st part: Pedestal and timing plots from BCal. 2nd part: higher level plots from BCal.
- 12/04/14: ST report (Eric Pooser) ST Performance Update
- 12/05/14: FDC report (Lubomir) FDC performance in fall run, Gain calibration
- 12/08/14: TagM: report (Richard Jones) Status Report on TAGM Commissioning
- 12/09/14: TagH: report (Nathan Sparks) Status Report on TAGH Commissioning
- 12/10/14: TagM: report (Richard Jones) Status Report on TAGM Commissioning, part 2
- 12/11/14: FCAL: report (Matt Shepherd) Pedestal Status Report
- 12/12/14: CDC: report (Mike) CDC Status Update
- 12/15/14: TOF report ( Alex. Ostrovidov) TOF low level signals
- 12/17/14: DAQ report (Sergei Furletov) DAQ report
- 12/18/14: Pair Spectrometer report (Alex. Somov)